Determine the power delivered by the turbine shaft if the steam at the inlet of the turbine is at 1MPa and 500 degrees. The steam at the outlet is saturated vapor and at 5KPa. The flow rate is 20kg/s.
1
Expert's answer
2021-07-21T09:05:49-0400

Given,

Pressure at the inlet turbine "(P_1)=1MPa=10^6 Pa"

Temperature "(T_1)=500^o C"

Outlet pressure "(P_2)=5KPa =5000Pa"

Flow rate "=20kg\/s"

"\eta =\frac{h_i -h_f}{h_i - h_f'}"


"=\frac{3479.1-2560.7}{3479.1-137.75}"


"=0.274"

Output energy "=m(h_1-h_o)"

"=20(3479.1-2560.7) =18368kw"

Output shaft "=18368\times \eta"

"=18368\times 0.274"

"=5032.05 kW"